# Self-embedding similitudes of Bedford-McMullen carpets with dependent ratios

### v2

@@ 3. The independent case
Let us pick a sequence $\{R_{k}\}_{k=1}^{\infty}$, where each $R_{k}$ is a level-$k$ rectangle and $R_{k+1}\subset R_{k}$ for all $k\geq 1$. Write $a_{k}$ to be the left bottom vertex of $R_{k}$ and $Q_{k}$ to be any $n^{-k}\times m^{-k}$ grid rectangle containing $f(a_{k})$. For convenience, we also let $\ell^{k},\ell_{k}$ be the bottom and left edges of $R_{k}$, respectively. So $\ell^{k}\cap\ell_{k}=\{a_{k}\}$. See Figure 2 for an illustration. Below we record two simple but key facts. Here we abuse notation slightly by writing $\varphi_{Q_{k}}$ to be the natural affine map sending $[0,1]^{2}$ onto $Q_{k}$.
@@
    $|\varphi_{Q_{k}}^{-1}f(\ell^{k})|$ $=\sqrt{n^{2k}|\pi_{1}(f(\ell^{k}))|^{2}+m^{2k}|\pi_{2}(f(\ell^{k}))|^{2}}$
    $=\sqrt{n^{2k}\cdot\lambda^{2}|\ell^{k}|^{2}\cdot\frac{1}{1+u^{2}}+m^{2k}\cdot\lambda^{2}|\ell^{k}|^{2}\cdot\frac{u^{2}}{1+u^{2}}}$
    $=\sqrt{n^{2k}\cdot\frac{\lambda^{2}}{n^{2k}}\cdot\frac{1}{1+u^{2}}+m^{2k}\cdot\frac{\lambda^{2}}{n^{2k}}\cdot\frac{u^{2}}{1+u^{2}}}$
    (3.2) $\to\frac{\lambda}{\sqrt{1+u^{2}}},\quad k\to\infty$
@@
    Similarly, the line supporting $\varphi_{Q_{k}}^{-1}f(\ell_{k})$ has slope $-\frac{m^{k}}{n^{k}u}$, which also tends to 00 as $k\to\infty$. Moreover, $\varphi_{Q_{k}}^{-1}f(\ell_{k})$ has length
    $|\varphi_{Q_{k}}^{-1}f(\ell_{k})|$ $=\sqrt{n^{2k}|\pi_{1}(f(\ell_{k}))|^{2}+m^{2k}|\pi_{2}(f(\ell_{k}))|^{2}}$
    $=\sqrt{n^{2k}\cdot\frac{\lambda^{2}}{m^{2k}}\cdot\frac{1}{1+u^{-2}}+m^{2k}\cdot\frac{\lambda^{2}}{m^{2k}}\cdot\frac{u^{-2}}{1+u^{-2}}}$
@@
Write $\widetilde{a}_{k}:=\varphi_{Q_{k}}^{-1}f(a_{k})$. Since $f(a_{k})\in Q_{k}$, $\widetilde{a}_{k}\in[0,1]^{2}$. Fix a large $k$ (will be specified later) so that $\frac{m^{k}}{n^{k}}\cdot\max\{u^{-1},u\}<\lambda$ and $|\varphi_{Q_{k}}^{-1}f(\ell^{k})|\leq\frac{2\lambda}{\sqrt{1+u^{2}}}$ (recall (3.2)). Let us truncate the parallelogram $\varphi_{Q_{k}}^{-1}f(R_{k})$ as $\bigcup_{t\in\mathbb{Z}}\widetilde{E}_{t}$, where
@@
(3.3) $=m^{k}\cdot\frac{\lambda}{n^{k}}\cdot\frac{u}{\sqrt{1+u^{2}}}+m^{k}\cdot\frac{\lambda}{m^{k}}\cdot\frac{u^{-1}}{\sqrt{1+u^{-2}}}<2\lambda.$
Let us prove Case 3 under the following Condition A first and then explain how to move on without its validity.
@@
There are integers $t_{0}\geq-1$ and $0\leq j_{0}\leq m$ and such that $|\pi_{1}(\widetilde{E}_{t_{0}})|=1$ while $\pi_{2}(\widetilde{E}_{t_{0}})\subset[\tfrac{j_{0}}{m},\tfrac{j_{0}+1}{m}]$.
Assume Condition A. Let $j^{\prime}_{0}:=j_{0}\pmod{m}$. Recall that $\varphi_{Q_{k}}^{-1}f(\ell_{k})$ is supported on a line of slope $-\frac{m^{k}}{n^{k}u}$, say $g(x)=-\frac{m^{k}}{n^{k}u}x+c$. Since #$\#I_{j^{\prime}_{0}}<n$, there is some $10\leq i_{0}\leq n-$ 1 such that $K$ does not intersect the open rectangle $(\frac{i_{0}}{n},\frac{i_{0}+1}{n})\times(\frac{j^{\prime}_{0}}{m},\frac{j^{\prime}_{0}+1}{m})=:U$. Write $e$ to be the point $(-t_{0},\lfloor\frac{j_{0}}{m}\rfloor)$ and $x_{0}:=-t_{0}+\frac{i_{0}+1/2}{n}$. Then the point $\xi_{0}:=(x_{0},g(x_{0}))$ lies on the segment $\varphi_{Q_{k}}^{-1}f(\ell_{k})$, and
@@
Since $\xi_{0}+\varphi_{Q_{k}}^{-1}f(\ell^{k})$ is a translated copy of $\varphi_{Q_{k}}^{-1}f(\ell^{k})$ and is contained in $\varphi_{Q_{k}}^{-1}f(R_{k})$ (the original flat parallelogram), $f^{-1}\varphi_{Q_{k}}(\xi_{0})+\ell^{k}$ is a translated copy of $\ell^{k}$ and is contained in $R_{k}$. Since $\ell^{k}$ is the bottom edge of the level-$k$ rectangle $R_{k}$, $\ell^{k}\cap\varphi_{R_{k}}(K)=\varphi_{R_{k}}(K^{0})$. So there exists some $y_{0}\in[0,1]$ such that
$(f^{-1}\varphi_{Q_{k}}(\xi_{0})+\ell^{k})\cap\varphi_{R_{k}}(K)=\varphi_{R_{k}}(K^{y_{0}})\neq\varnothing,$
@@
(3.5) $(\xi_{0}+\varphi_{Q_{k}}^{-1}f(\ell^{k}))\cap\varphi_{Q_{k}}^{-1}f\varphi_{R_{k}}(K)=\varphi_{Q_{k}}^{-1}f\varphi_{R_{k}}(K^{y_{0}})\neq\varnothing.$
@@
$\subset(U+e)\cap\Big{(}\bigcup_{\begin{subarray}{c}\text{$R$ is alevel-$k$ rectangle}\\ \text{$R\cap f(K)\neq\varnothing$}\end{subarray}}\varphi_{Q_{k}}^{-1}\varphi_{R}(K)\Big{)}$ (since $f(K)\subset K$)
@@
Now, suppose Condition A is not fulfilled. Set $j_{*}:=\min\{j\in\mathbb{Z}:\pi_{2}(\widetilde{a}_{k})<\frac{j}{m}\}\geq 1$. So the segment $[-2,-1]\times\{\frac{j_{*}}{m}\}$ should intersect $\widetilde{E}_{2}$: otherwise, Condition A holds for $\widetilde{E}_{2}$ and $j_{*}$ or $j_{*}-1$. We claim that
(3.6) $\min\pi_{2}(\widetilde{E}_{2}\cap(\{-2\}\times\mathbb{R}))<\tfrac{j_{*}}{m}<\max\pi_{2}(\widetilde{E}_{2}\cap(\{-1\}\times\mathbb{R})).$
Otherwise, either $\max\pi_{2}(\widetilde{E}_{2}\cap(\{-1\}\times\mathbb{R}))\leq\frac{j_{*}}{m}$ and thus Condition A holds for $\widetilde{E}_{1}$ and $j_{*}-1$; or $\min\pi_{2}(\widetilde{E}_{2}\cap(\{-2\}\times\mathbb{R}))\geq\frac{j_{*}}{m}$ and thus Condition A holds for $\widetilde{E}_{3}$ and $j_{*}$.
If #$\#I_{j^{\prime}}\geq 2$ for some $j^{\prime}$, then we can pick $R_{k}$ properly at the beginning such that
$R_{k}=\varphi_{i_{1},j^{\prime}}\varphi_{i_{2},j^{\prime}}\cdots\varphi_{i_{k},j^{\prime}}([0,1]^{2}),$
where $i_{1},\ldots,i_{k}\in I_{j^{\prime}}$. As a result, selecting $i^{\prime}_{k}\in I_{j^{\prime}}\setminus\{i_{k}\}$, $R^{\prime}_{k}:=\varphi_{i_{1},j^{\prime}}\cdots\varphi_{i_{k-1},j^{\prime}}\varphi_{i^{\prime}_{k},j^{\prime}}([0,1]^{2})$ is another level-$k$ rectangle such that $\pi_{2}(R^{\prime}_{k})=\pi_{2}(R_{k})$ (i.e, they are on the same row of the $n^{-k}\times m^{-k}$ grid) while $\pi_{1}(R^{\prime}_{k})$ is a $\frac{|i^{\prime}_{k}-i_{k}|}{n^{k}}$-translation of $\pi_{1}(R_{k})$. Note that $\varphi_{Q_{k}}^{-1}f(R^{\prime}_{k})$ is simply a translated copy of $\varphi_{Q_{k}}^{-1}f(R_{k})$ along the line supporting $\varphi_{Q_{k}}^{-1}f(\ell^{k})$, and they have disjoint interiors. Picking $k$ large at the beginning (as long as $|\pi_{1}(\widetilde{E}_{2})|=|\pi_{1}(\widetilde{E}^{\prime}_{2})|=1$), the set
$\widetilde{E}^{\prime}_{2}:=\varphi_{Q_{k}}^{-1}f(R^{\prime}_{k})\cap\{x\in\mathbb{R}^{2}:\pi_{1}(x)\in[-2,-1]\},$
is simply a translated copy of $\widetilde{E}_{2}$ (see Figure 4 for an illustration). By (3.6), the segment $[-2,-1]\times\{\frac{j_{*}}{m}\}$ lies entirely in $\widetilde{E}_{2}$ and hence does not meet the interior of $\widetilde{E}^{\prime}_{2}$. Moreover, by the definition of $R^{\prime}_{k}$, it is not hard to check that either (when $\widetilde{E}^{\prime}_{2}$ lies above that segment)
$\frac{j_{*}}{m}\leq\min\pi_{2}(\widetilde{E}^{\prime}_{2})<\max\pi_{2}(\widetilde{E}^{\prime}_{2})\leq\min\pi_{2}(\widetilde{E}_{2})+2n\lambda<\frac{j_{*}+1}{m};$
or (when $\widetilde{E}^{\prime}_{2}$ lies below that segment)
$\frac{j_{*}}{m}\geq\max\pi_{2}(\widetilde{E}^{\prime}_{2})>\min\pi_{2}(\widetilde{E}^{\prime}_{2})\geq\frac{j_{*}}{m}-2n\lambda>\frac{j_{*}-1}{m}.$
Without loss of generality, let us assume the former. So Condition A holds for $\widetilde{E}^{\prime}_{2}$ and $j_{*}$. Applying the previous argument (i.e., the one after assuming Condition A), one can show that the parallelogram $\varphi_{Q_{k}}^{-1}f(R^{\prime}_{k})$ passes through a hole in $\bigcup_{z\in\mathbb{Z}^{2}}K+z$, which leads to a contradiction as before.
[…]
Figure 4. An illustration of the possible positions of $\widetilde{E}^{\prime}_{2}$ (color online)
Finally, if #$\#I_{j}=1$ for all $j$, then since $n>m\geq 2$, we can find $10\leq i_{*}\leq n-$ 1 such that $(i_{*},j_{*}\pmod{m}),(i_{*},j_{*}-1)\notin\Lambda$. Write $j^{\prime}_{*}:=j_{*}\pmod{m}$. Then, since $\pi_{2}(\widetilde{E}_{2})\subset[\tfrac{j_{*}-1}{m},\tfrac{j_{*}+1}{m}]$, applying the previous argument (i.e., the one after assuming Condition A) to $t_{0}=2$, $j_{0}=j_{*}$, $i_{0}=i_{*}$, $j^{\prime}_{0}=j^{\prime}_{*}$ and $U=(\frac{i_{0}}{n},\frac{i_{0}+1}{n})\times(\frac{j^{\prime}_{*}-1}{m},\frac{j^{\prime}_{*}+1}{m})$, one can show that the parallelogram $\varphi_{Q_{k}}^{-1}f(R_{k})$ passes through a hole (which is a translatied copy of $U$) in $\bigcup_{z\in\mathbb{Z}^{2}}K+z$, which leads to a contradiction as before. ∎
@@ 4. The dependent self-affine case
In this section, we fix a Bedford-McMullen carpet $K=K(n,m,\Lambda)$ that is not supported in any line and assume that $\frac{\log n}{\log m}\in\mathbb{Q}$. Without loss of generality, write $n=m^{p/q}$, where $p,q\in\mathbb{Z}^{+}$ are coprime. Since $n>m$, $p>q$. Recall that $N=\max_{j}\#I_{j}$.
@@
Let $f\in\mathcal{S}(\mathbb{R}^{2})$. If $f(K)\subset K$, then we can find $g\in\mathcal{S}(R^{2})$ and a compact set $C\subset\mathbb{R}$ with $\operatorname{dim_{H}}C\leq\frac{\log N}{\log n}$ such that $g(K)\subset C\times\mathbb{R}$. Moreover, if $f$ is oblique, then $g$ can be picked oblique as well.
@@
Since the contraction ratio of $f^{t}$ is $\lambda^{t}$, $K\subset[0,1]^{2}$ and $f^{t}(K)\subset K$, $f^{t}(K)$ is contained in at most $9$ squares of side length $m^{-pk_{t}}=n^{-qk_{t}}$. More precisely, there are $x\in\frac{\mathbb{N}}{n^{qk_{t}}},y\in\frac{\mathbb{N}}{m^{pk_{t}}}$ such that
$f^{t}(K)\subset(x,x+3n^{-qk_{t}})\times(y,y+3m^{-pk_{t}})=:Q_{t}.$
@@
Moreover, write $y=\sum_{k=1}^{pk_{t}}y_{k}m^{-k}$, where $y_{k}\in\{0,\ldots,m-1\}$. Since $p>q$, the collection of all level-$pk_{t}$ rectangles in $\overline{Q^{\prime}_{t}}$, denoted by $\mathcal{R}_{t}$, is (if non-empty)
@@
$\subset\bigcup_{\begin{subarray}{c}(x_{k},y_{k})\in\Lambda,\\ qk_{t}+1\leq k\leq pk_{t}\end{subarray}}\Big{[}\sum_{k=qk_{t}+1}^{pk_{t}}\frac{x_{k}}{n^{k-qk_{t}}},n^{-(p-q)k_{t}}+\sum_{k=qk_{t}+1}^{k_{t}}\frac{x_{k}}{n^{k-qk_{t}}}\Big{]}\times\mathbb{R}$
@@
Suppose $g(K)\cap(0,1)^{2}\neq\varnothing$. Then $h_{t}^{-1}f^{t}(K)\cap(0,1)^{2}\neq\varnothing$ for all large $t$. For each of these $t$, note that $\mathcal{R}_{t}\neq\varnothing$ (recall (4.2)) and write $\bm{j}_{t}:=y_{qk_{t}+1}\cdots y_{pk_{t}}$, which is a word of length $(p-q)k_{t}\to\infty$ as $t\to\infty$. Using Cantor’s diagonal argument, it is not hard to find an infinite word $\bm{j}=j_{1}j_{2}\cdots\in J^{\infty}$ such that for all $M\geq 1$, there is a large $t$ such that $|\bm{j}_{t}\wedge\bm{j}|\geq M$, where $\bm{j}_{t}\wedge\bm{j}$ denotes their longest common prefix and ||$\cdot$|| ⋅ | stands for the length. Let $z:=\sum_{k=1}^{\infty}j_{k}m^{-k}$.
@@
If not, then there is some $b\in g(K)\cap(0,1)^{2}$ such that
$\delta:=\operatorname{dist}(\{b\},K^{z}\times\mathbb{R})>0.$
where $\operatorname{dist}(A,B):=\inf\{|a-b|:a\in A,b\in B\}$ is the usual distance. Let $M$ be a large integer with $n^{-M}<\frac{\delta}{4}$. Pick $t$ so large that the following conditions hold:
@@
    $d_{H}(g(K)\cap(0,1)^{2},h_{t}^{-1}f^{t}(K)\cap(0,1)^{2})<\frac{\delta}{4}$ (where $d_{H}$ is the Hausdorff distance, see [8]);
@@
The first condition gives us a $b_{1}\in h_{t}^{-1}f^{t}(K)\cap(0,1)^{2}$ with $|b_{1}-b|<\frac{\delta}{4}$. Since $|\bm{j}_{t}\wedge\bm{j}|\geq M$, it is not hard to see from the definition of $E_{t}$ and $\bm{j}_{t}$ that $\operatorname{dist}(E_{t},K^{z})\leq n^{-M}$. Hence
$\operatorname{dist}(\{b_{1}\},K^{z}\times\mathbb{R})$ $\leq\operatorname{dist}(h_{t}^{-1}f^{t}(K)\cap(0,1)^{2},K^{z}\times\mathbb{R})$
$\leq\operatorname{dist}(E_{t},K^{z})$ (by (4.3))
$\leq n^{-M}<\delta/4.$
But it follows that
$\delta=\operatorname{dist}(\{b\},K^{z}\times\mathbb{R})\leq|b-b_{1}|+\operatorname{dist}(\{b_{1}\},K^{z}\times\mathbb{R})<\delta/2,$
which is impossible.
Applying the same argument to the other eight squares in $Q_{t}$, we find at most nine points $z_{0,0}:=z,z_{0,1},\ldots,z_{2,2}\in\pi_{2}(K)$ such that
@@
We also remark that the integer $k_{t}$ defined in (4.1) equals $\lfloor(-\frac{t\log\lambda}{\log m}-1)/p\rfloor$. So if $\frac{\log\lambda}{\log n}\notin\mathbb{Q}$, then $\frac{\log\lambda}{\log m}\notin\mathbb{Q}$ and hence $\{n^{qk_{t}}\lambda^{t}\}_{t}$ is dense in an open subinterval of $(0,1)$. Since the proof of Lemma 4.1 works for all convergent subsequence of $\{h_{t}^{-1}f^{t}\}_{t=1}^{\infty}$, one can find such an accumulation map $\widetilde{g}$ so that $\widetilde{g}(K)\subset C_{\widetilde{g}}\times\mathbb{R}$, where $C_{\widetilde{g}}$ is a closed set with $\operatorname{dim_{H}}C\leq\frac{\log N}{\log n}$ (just as $C$ in Lemma 4.1), and the contraction ratio of $\widetilde{g}$ is an irrational power of $n$.
Now we are able to prove the non-obliqueness theorem for the dependent case.

### v3

@@ 3. The independent case
Let us pick a sequence $\{R_{k}\}_{k=1}^{\infty}$, where each $R_{k}$ is a level-$k$ rectangle and $R_{k+1}\subset R_{k}$ for all $k\geq 1$. Write $a_{k}$ to be the left bottom vertex of $R_{k}$ and $Q_{k}$ to be any $n^{-k+p}\times m^{-k+p}$ grid rectangle containing $f(a_{k})$. For convenience, we also let $\ell^{k},\ell_{k}$ be the bottom and left edges of $R_{k}$, respectively. So $\ell^{k}\cap\ell_{k}=\{a_{k}\}$. See Figure 2 for an illustration. Below we record two simple but key facts. Here we abuse notation slightly by writing $\varphi_{Q_{k}}$ to be the natural affine map sending $[0,1]^{2}$ onto $Q_{k}$.
@@
    $|\varphi_{Q_{k}}^{-1}f(\ell^{k})|$ $=\sqrt{n^{2(k-p)}|\pi_{1}(f(\ell^{k}))|^{2}+m^{2(k-p)}|\pi_{2}(f(\ell^{k}))|^{2}}$
    $=\sqrt{n^{2(k-p)}\cdot\lambda^{2}|\ell^{k}|^{2}\cdot\frac{1}{1+u^{2}}+m^{2(k-p)}\cdot\lambda^{2}|\ell^{k}|^{2}\cdot\frac{u^{2}}{1+u^{2}}}$
    $=\sqrt{n^{2(k-p)}\cdot\frac{\lambda^{2}}{n^{2k}}\cdot\frac{1}{1+u^{2}}+m^{2(k-p)}\cdot\frac{\lambda^{2}}{n^{2k}}\cdot\frac{u^{2}}{1+u^{2}}}$
    (3.2) $\to\frac{n^{-p}\lambda}{\sqrt{1+u^{2}}},\quad k\to\infty$
@@
    Similarly, the line supporting $\varphi_{Q_{k}}^{-1}f(\ell_{k})$ has slope $-\frac{m^{k-p}}{n^{k-p}u}$, which also tends to 00 as $k\to\infty$. Moreover, $\varphi_{Q_{k}}^{-1}f(\ell_{k})$ has length
    $|\varphi_{Q_{k}}^{-1}f(\ell_{k})|$ $=\sqrt{n^{2(k-p)}|\pi_{1}(f(\ell_{k}))|^{2}+m^{2(k-p)}|\pi_{2}(f(\ell_{k}))|^{2}}$
    $=\sqrt{n^{2(k-p)}\cdot\frac{\lambda^{2}}{m^{2k}}\cdot\frac{1}{1+u^{-2}}+m^{2(k-p)}\cdot\frac{\lambda^{2}}{m^{2k}}\cdot\frac{u^{-2}}{1+u^{-2}}}$
@@
Write $\widetilde{a}_{k}:=\varphi_{Q_{k}}^{-1}f(a_{k})$. Since $f(a_{k})\in Q_{k}$, $\widetilde{a}_{k}\in[0,1]^{2}$. Fix a large $k$ (will be specified later) so that $\frac{m^{k-p}}{n^{k-p}}\cdot\max\{u^{-1},u\}<\lambda$ and $|\varphi_{Q_{k}}^{-1}f(\ell^{k})|\leq\frac{2n^{-p}\lambda}{\sqrt{1+u^{2}}}$ (recall (3.2)). Let us truncate the parallelogram $\varphi_{Q_{k}}^{-1}f(R_{k})$ as $\bigcup_{t\in\mathbb{Z}}\widetilde{E}_{t}$, where
@@
(3.3) $=m^{k-p}\cdot\frac{\lambda}{n^{k}}\cdot\frac{u}{\sqrt{1+u^{2}}}+m^{k-p}\cdot\frac{\lambda}{m^{k}}\cdot\frac{u^{-1}}{\sqrt{1+u^{-2}}}<2m^{-p}\lambda.$
Let us prove Case 3 under the following Condition A first and then show its validity.
@@
There are integers $t_{0}\geq-1$ and $0\leq j_{0}\leq m$ such that $|\pi_{1}(\widetilde{E}_{t_{0}})|=1$ while $\pi_{2}(\widetilde{E}_{t_{0}})\subset[\tfrac{j_{0}}{m},\tfrac{j_{0}+1}{m}]$.
Assume that Condition A holds. Let $j^{\prime}_{0}:=j_{0}\pmod{m}$. Recall that $\varphi_{Q_{k}}^{-1}f(\ell_{k})$ is supported on a line of slope $-\frac{m^{k-p}}{n^{k-p}u}$, say $g(x)=-\frac{m^{k-p}}{n^{k-p}u}x+c$ (the equation of that line). Since #$\#I_{j^{\prime}_{0}}<n$, there is some $10\leq i_{0}\leq n-$ 1 such that $K$ does not intersect the open rectangle $(\frac{i_{0}}{n},\frac{i_{0}+1}{n})\times(\frac{j^{\prime}_{0}}{m},\frac{j^{\prime}_{0}+1}{m})=:U$. Write $e$ to be the point $(-t_{0},\lfloor\frac{j_{0}}{m}\rfloor)$ and $x_{0}:=-t_{0}+\frac{i_{0}+1/2}{n}$. Then the point $\xi_{0}:=(x_{0},g(x_{0}))$ lies on the segment $\varphi_{Q_{k}}^{-1}f(\ell_{k})$, and
@@
Since $\xi_{0}+\varphi_{Q_{k}}^{-1}f(\ell^{k})$ is a translated copy of $\varphi_{Q_{k}}^{-1}f(\ell^{k})$ and is contained in $\varphi_{Q_{k}}^{-1}f(R_{k})$, $f^{-1}\varphi_{Q_{k}}(\xi_{0})+\ell^{k}$ is a translated copy of $\ell^{k}$ and is contained in $R_{k}$. Since $\ell^{k}$ is the bottom edge of the level-$k$ rectangle $R_{k}$, $\ell^{k}\cap\varphi_{R_{k}}(K)=\varphi_{R_{k}}(K^{0})$. So there is $y_{0}\in[0,1]$ such that
$(f^{-1}\varphi_{Q_{k}}(\xi_{0})+\ell^{k})\cap\varphi_{R_{k}}(K)=\varphi_{R_{k}}(K^{y_{0}}\times\{y_{0}\})\neq\varnothing,$
@@
(3.5) $(\xi_{0}+\varphi_{Q_{k}}^{-1}f(\ell^{k}))\cap\varphi_{Q_{k}}^{-1}f\varphi_{R_{k}}(K)=\varphi_{Q_{k}}^{-1}f\varphi_{R_{k}}(K^{y_{0}}\times\{y_{0}\})\neq\varnothing.$
@@
$\subset(U+e)\cap\Big{(}\bigcup_{\begin{subarray}{c}\text{$R$ is alevel-$(k-p)$ rectangle}\\ \text{$R\cap f(K)\neq\varnothing$}\end{subarray}}\varphi_{Q_{k}}^{-1}\varphi_{R}(K)\Big{)}$ (since $f(K)\subset K$)
@@
It remains to prove Condition A. The key is to note that for every $y$, $\varphi_{Q_{k}}^{-1}f(R_{k})\cap(\mathbb{R}\times\{y\})$ (which is a horizontal slice of the parallelogram) has length at most $Cn^{-p}\lambda$. To see this, just look at Figure 4 and recall that the two acute edges of that triangle has slope $\frac{m^{k-p}u}{n^{k-p}}$ and $-\frac{m^{k-p}}{n^{k-p}u}$, respectively.
$\widetilde{a}_{k}$$\varphi_{Q_{k}}^{-1}f(\ell^{k})$
Figure 4. Length of horizontal slices of $\varphi_{Q_{k}}^{-1}f(R_{k})$
If Condition A holds for $t_{0}=1$ and some suitable $j_{0}$ then we are done. Otherwise, there exists some $j_{*}$ such that $\widetilde{E}_{1}\cap(\mathbb{R}\times\{\frac{j_{*}}{m}\})\neq\varnothing$. Since $|\varphi_{Q_{k}}^{-1}f(R_{k})\cap(\mathbb{R}\times\{\frac{j_{*}}{m}\})|\leq Cn^{-p}\lambda$, picking $k$ large at the beginning, it is not hard to check that Condition A holds for $t_{*}$ and $j_{*}$, where $t_{*}:=\min\{t\geq 2:\widetilde{E}_{t}\cap(\mathbb{R}\times\{\tfrac{j_{*}}{m}\})=\varnothing\}$. ∎
@@ 4. The dependent self-affine case
In this section, we fix a Bedford-McMullen carpet $K=K(n,m,\Lambda)$ that is not supported in any line and assume that $\frac{\log n}{\log m}\in\mathbb{Q}$. Without loss of generality, write $n=m^{p/q}$, where $p,q\in\mathbb{Z}^{+}$ are coprime. It follows from $n>m$ that $p>q$. Recall that $N=\max_{j}\#I_{j}$.
@@
Let $f\in\mathcal{S}(\mathbb{R}^{2})$ be a contracting map. If $f(K)\subset K$, then we can find $g\in\mathcal{S}(\mathbb{R}^{2})$ and a compact set $C\subset\mathbb{R}$ with $\operatorname{dim_{H}}C\leq\frac{\log N}{\log n}$ such that $g(K)\subset C\times\mathbb{R}$. Moreover, if $f$ is oblique, then $g$ can be picked oblique as well.
@@
Since the contraction ratio of $f^{t}$ is $\lambda^{t}$, $K\subset[0,1]^{2}$ and $f^{t}(K)\subset K$, $f^{t}(K)$ is contained in at most $4$ squares of side length $m^{-pk_{t}}=n^{-qk_{t}}$. More precisely, there are $x\in n^{-qk_{t}}\mathbb{N},y\in m^{-pk_{t}}\mathbb{N}$ such that
$f^{t}(K)\subset(x,x+2n^{-qk_{t}})\times(y,y+2m^{-pk_{t}})=:Q_{t}.$
@@
Suppose $g(K)\cap(0,1)^{2}\neq\varnothing$. Then $h_{t}^{-1}f^{t}(K)\cap(0,1)^{2}\neq\varnothing$ for all large $t$. For each of these $t$, write $y=\sum_{k=1}^{pk_{t}}y_{k}m^{-k}$, where $y_{k}\in\{0,\ldots,m-1\}$. Since $p>q$, the collection of all level-$pk_{t}$ rectangles in $\overline{Q^{\prime}_{t}}$, denoted by $\mathcal{R}_{t}$, is
@@
$\subset\bigcup_{\begin{subarray}{c}(x_{k},y_{k})\in\Lambda,\\ qk_{t}+1\leq k\leq pk_{t}\end{subarray}}\Big{[}\sum_{k=qk_{t}+1}^{pk_{t}}\frac{x_{k}}{n^{k-qk_{t}}},n^{-(p-q)k_{t}}+\sum_{k=qk_{t}+1}^{pk_{t}}\frac{x_{k}}{n^{k-qk_{t}}}\Big{]}\times\mathbb{R}$
@@
Moreover, write $\bm{j}_{t}:=y_{qk_{t}+1}\cdots y_{pk_{t}}$, which is a word of length $(p-q)k_{t}\to\infty$ as $t\to\infty$. Using Cantor’s diagonal argument, it is not hard to find an infinite word $\bm{j}=j_{1}j_{2}\cdots\in J^{\infty}$ such that for all $M\geq 1$, there is a large $t$ such that $|\bm{j}_{t}\wedge\bm{j}|\geq M$, where $\bm{j}_{t}\wedge\bm{j}$ denotes their longest common prefix and ||$\cdot$|| ⋅ | stands for the length. Let $z:=\sum_{k=1}^{\infty}j_{k}m^{-k}$.
@@
Fix any $\delta>0$. Let $M$ be a large integer with $n^{-M}<\frac{\delta}{2}$. Pick $t$ so large that the following conditions hold:
@@
    $g(K)\cap(0,1)^{2}\subset\mathcal{N}_{\delta/2}(h_{t}^{-1}f^{t}(K)\cap(0,1)^{2})$, where $\mathcal{N}_{\delta/2}(\cdot)$ denotes the $\delta/2$-neighborhood.
@@
Since $|\bm{j}_{t}\wedge\bm{j}|\geq M$, it is not hard to see from the definition of $E_{t}$ and $\bm{j}_{t}$ that $E_{t}\subset\mathcal{N}_{n^{-M}}(K^{z})$. Hence
$g(K)\cap(0,1)^{2}$ $\subset\mathcal{N}_{\delta/2}(h_{t}^{-1}f^{t}(K)\cap(0,1)^{2})$ (condition (1))
$\subset\mathcal{N}_{\delta/2}(E_{t}\times\mathbb{R})$ (by (4.3))
$\subset\mathcal{N}_{n^{-M}+\delta/2}(K^{z}\times\mathbb{R})\subset\mathcal{N}_{\delta}(K^{z}\times\mathbb{R}).$
Since $\delta$ is arbitrary, we establish (4.4).
Applying the same argument to the other three squares in $Q_{t}$, we find at most four points $z_{0,0}:=z,z_{0,1},z_{1,0},z_{1,1}\in\pi_{2}(K)$ such that
@@
We also remark that the integer $k_{t}$ defined in (4.1) equals $\lfloor(-\frac{t\log\lambda}{\log m}-1)/p\rfloor$. So if $\frac{\log\lambda}{\log n}\notin\mathbb{Q}$, then $\frac{\log\lambda}{\log m}\notin\mathbb{Q}$ and hence $\{n^{qk_{t}}\lambda^{t}\}_{t}$ is dense in an open subinterval of $(0,1)$. Since the proof of Lemma 4.1 works for all convergent subsequence of $\{h_{t}^{-1}f^{t}\}_{t=1}^{\infty}$, one can find such an accumulation map $\widetilde{g}$ so that $\widetilde{g}(K)\subset C_{\widetilde{g}}\times\mathbb{R}$, where $C_{\widetilde{g}}$ is a closed set with $\operatorname{dim_{H}}C_{\widetilde{g}}\leq\frac{\log N}{\log n}$ (just as $C$ in Lemma 4.1), and the contraction ratio of $\widetilde{g}$ is an irrational power of $n$.
Now we are able to prove the non-obliqueness theorem for the dependent case. The proof is separated into two parts, depending on whether $f$ is a contracting map or an isometry.